If a spring of stiffness ‘k’ is cut into parts ‘A’ and ‘B’ of length lA : lB = 2 : 3, then the stiffness of spring ‘A’ is given by
(a) 3k/5
(b) 2k/5
(c) k
(d) 5/2k

Grade:11

1 Answers

Kevin Nash
9 years ago

(d) Here, lA = 2l/5, lB = (3l/5)

kl = kAlA = kBlB kl = kA(2l/5

kA = 5k/2 ? kB =5k/3